OpenStudy (anonymous):

standard form 5 + 3i + 4 - 7i?

OpenStudy (anonymous):

9 - 4i

OpenStudy (zzr0ck3r):

standard form is a+bi

OpenStudy (zzr0ck3r):

you should try and lead people to the answer, not give them the answer straight off the bat @gaara438125

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

plz explain

OpenStudy (anonymous):

you should not tell me or correct me without trying to be polite about it, and not just demand i do something about it, thus leading me to disrespect you so there...how shiv asked " plz explain" now i explain @zzr0ck3r

OpenStudy (zzr0ck3r):

Read the rules @gaara438125 this has nothing to do with what I think

OpenStudy (anonymous):

5 + 4 = 9 3i - 7i = -4i <-- this has a variable which is a number we don't know thus making it so we cannot add it to the 9

OpenStudy (zzr0ck3r):

Im sure if you reread what I wrote you will see that I was in no way not polite...

OpenStudy (zzr0ck3r):

its not a variable we dont know... i = sqrt(-1) i.e not unknown.

OpenStudy (anonymous):

i does not always equal an imaginary number -.- plus the sqrt of a negative number can also be known as A or B or Elephant...it's a variable for a reason

OpenStudy (zzr0ck3r):

if they were not talking about sqrt(-1) there would be no "standard" form... so of course we are talking about complexe numbers. look up meta variables vs variables then come back and sprout some more stuff

OpenStudy (anonymous):

i'm not sure what a complexe number is and complex numbers don't have to deal with imaginary numbers...

OpenStudy (anonymous):

sorry, complex numbers do yes, but not to put something in standard form.

OpenStudy (zzr0ck3r):

what ?? every real number is a complex number of the standard form a + 0i where a is an element in R.

OpenStudy (anonymous):

only if a number of the form a+bi where a and b are real numbers and i is the square root of -1. not every real number...

OpenStudy (zzr0ck3r):

every real number...google man. the real numbers are a subset of the complex numbers. google then tell me sorry that you are wrong again:0

OpenStudy (anonymous):

leave it

OpenStudy (anonymous):

(complex number) (mathematics) a number of the form a+bi where a and b are real numbers and i is the square root of -1. would this definition FROM your source google help you?

OpenStudy (zzr0ck3r):

I will now teach you to google. google "is every real number a complex number" without the "" of course. then get back to me.

OpenStudy (anonymous):

In the special case where b=0...did you read? lolol shiv now i'm done i'm sure your getting annoyed about two stubborn morons arguing online

OpenStudy (zzr0ck3r):

every rea number is a special case of b = 0

OpenStudy (zzr0ck3r):

3+0i = 3 natural numbers are a subset of the integers integers are a subset of real numbers real numbers are a subset of the complex numbers when A is a subset of B then every element in A is in B. thus every real numbers is a complexe numbers by definition...